OpenStudy (anonymous):

What does x= when (-1,600,000/x^2)+.1=0? Thanks for your help!

OpenStudy (anonymous):

huh please retype your question by clicking the draw option below! because i cant seem to understand it

OpenStudy (campbell_st):

well the equation looks like \[\frac{-1600000}{x^2} + 0.1 = 0\] start by subtracting 0.1 from both sides of the equation the multiplying both sides by x^2 and you get \[-1600000 = -0.1x^2\] multiply both sides of the equation by -10 and you get \[16000000 = x^2\] now you can solve for x